OpenStudy (anonymous):

Find the slope of the line passing through the points (3, 4) and (2, 1).

OpenStudy (anonymous):

slope of line passing through points(x1,y1) and (x2,y2) is equal to slope= (y2-y1)/(x2-x1) = (-3)/(-1) =3